OpenStudy (anonymous):

If f(x) = 3x2 - x, find f(-2).

OpenStudy (solomonzelman):

\(\large\color{royalblue}{ f(x) = 3x^2 - x }\). If I was to find \(\large\color{red}{ f(-4) }\) I would plug in 4 for x, like this: \(\large\color{royalblue}{ f(\color{red}{-4}) = 3(\color{red}{-4})^2 - (\color{red}{-4})}\) \(\large\color{royalblue}{ f(\color{red}{-4}) = 3(16) +4}\) \(\large\color{royalblue}{ f(\color{red}{-4}) = 48+4}\) \(\large\color{royalblue}{ f(\color{red}{-4}) = 52}\)

OpenStudy (solomonzelman):

find \(\large\color{red}{ f(-4) }\) the same exact way, but you are plugging not -4, but -2 for x.

OpenStudy (anonymous):

the answer = 14?

OpenStudy (solomonzelman):

oh, typo in a second line, first reply. I would plug in `-` 4 for x.

OpenStudy (solomonzelman):

had too many typos, I deleted my last reply, but \(\large\color{royalblue}{ f(\color{red}{-2}) = 14}\) is correct.

OpenStudy (solomonzelman):

\(\large\color{royalblue}{ f(\color{red}{x}) = 3(x)^2-x}\) \(\large\color{royalblue}{ f(\color{rd}{-2}) = 3(-2)^2-(-2)}\) \(\large\color{royalblue}{ f(\color{rd}{-2}) = 3(4)-(-2)}\) \(\large\color{royalblue}{ f(\color{rd}{-2}) = 12+2}\) \(\large\color{royalblue}{ f(\color{red}{-2}) = 14}\)